# GREAT: Geometry-Intention Collaborative Inference for Open-Vocabulary 3D Object Affordance Grounding

### v1

@@ 5.1 Benchmark Setting
Evalization Metrics. For a thorough assessment, we use previous evaluation metrics from advanced works [56, 27] on 3D object affordance grounding to benchmark the model on our PIADv2 dataset, which include Area Under Curve [31] , average Intersection Over Union [45], SIMilarity [47], Mean Absolute Error [52].
@@
Implementation Details. We take PointNet++ [44] and ResNet18 [9] as the default 3D and 2D backbones, respectively. To ensure a fair comparison, the same feature extractor and settings are used to reproduce the baselines. We train the GREAT for 65 epochs with a batch size of 16, utilizing the Adam optimizer with a learning rate of 1e-4. More details about the benchmark setting can be found in the appendix.
@@ 5.4 Performance Analysis
Multiple Affordances. To verify whether the model reasons about the 3D object affordance regions based on the understanding of interaction images, we use the model to infer distinct affordances for the same objects, as shown in Fig. 7. The results demonstrate that, for the same object, the model outputs different results depending on the interaction, and the localized 3D object affordance regions are consistent with the interactions depicted in the 2D images. This indicates that the model reasons about the interaction details and localizes the affordance regions based on the interaction images, rather than relying solely on a direct category-driven mapping.
@@
Figure 8: Multiple Instances. (a) Similar geometrics instances. (b) Different geometric instances.
@@ 6 Conclusion
We present grounding 3D object affordance in an open-vocabulary fashion, which reasons from interaction images, extrapolating from predefined sample space and generalize to unseen scenarios. To achieve so, we propose a novel framework to utilize multi-head affordance chain of thought reasoning, excavating invariant geometric properties and analogous reasoning about potential interactions, with the alignment of cross-modal features to localize 3D object affordance region. Furthermore, We introduce the largest 3D object affordance dataset PIADv2, which contains 15K interaction images and over 38K 3D objects with annotations. Extensive experiments demonstrate the effectiveness and superiority of GREAT. It supports affordance understanding in open scenarios, potentially enhancing robots’ autonomous interaction in unknown environments. We believe it could offer fresh insights and promote research in the area of visual affordance understanding.

### v2

@@ 5.1 Benchmark Setting
Evaluation Metrics. For a thorough assessment, we use previous evaluation metrics from advanced works [56, 25] to benchmark the model on our PIADv2 dataset, which include Area Under Curve [29] , average Intersection Over Union [45], SIMilarity [47], Mean Absolute Error [52].
@@
Implementation Details. We take PointNet++ [43] and ResNet18 [9] as the default 3D and 2D backbones, respectively. To ensure a fair comparison, the same feature extractor and settings are used to reproduce the baselines. We train the GREAT for 65 epochs with a batch size of 16, utilizing the Adam optimizer with a learning rate of 1e-4.
@@ 5.4 Performance Analysis
Multiple Affordances. To verify whether the model reasons about the 3D object affordance regions based on the understanding of interaction images, we use the model to infer distinct affordances for the same objects, as shown in Fig. 7. The results demonstrate that, for the same object, the model outputs different results depending on the interaction, and the localized 3D object affordance regions are consistent with the interactions depicted in the 2D images.
@@
Figure 8: Multiple Instances. (a) Similar geometric instances. (b) Different geometric instances.
@@ 6 Conclusion
We present grounding 3D object affordance in an open-vocabulary fashion, which reasons from interaction images, extrapolating from predefined sample space and generalize to unseen scenarios. To achieve so, we propose a novel framework to utilize multi-head affordance chain-of-thought reasoning, excavating invariant geometric properties and analogous reasoning about potential interactions, with the alignment of cross-modal features to localize 3D object affordance region. Furthermore, We introduce the largest 3D object affordance dataset PIADv2, which contains 15K interaction images and over 38K 3D objects with annotations. Extensive experiments demonstrate the effectiveness and superiority of GREAT. It supports affordance understanding in open scenarios, potentially enhancing robots’ autonomous interaction in unknown environments. We believe it could offer fresh insights and promote research in the area of visual affordance understanding.
@@
Acknowledgments. This work is supported by the National Natural Science Foundation of China (NSFC) under Grants 62306295, 62225207 and 62436008.